OpenStudy (goformit100):

How many words can be formed with the letter of the word 'VICE-CHANCELLOR' so that the vowels are together ?

OpenStudy (goformit100):

@mathslover

OpenStudy (shamim):

here vowels r i,a,e,o

OpenStudy (shamim):

there 5 vowels

OpenStudy (shamim):

E=2

OpenStudy (shamim):

consonant=9

OpenStudy (shamim):

u can think all the vowels r just only 1 letter

OpenStudy (shamim):

so 10 letter can make different word by\[\frac{ 10! }{ 3!2! }\]

OpenStudy (shamim):

i divided by\[3!\] and \[2!\]

OpenStudy (shamim):

because there r 3 C and 2 L

OpenStudy (shamim):

and we can arrange 5 vowels by\[\frac{ 5! }{ 2! }\]

OpenStudy (shamim):

there 2 E so divided by\[2!\]

OpenStudy (shamim):

so the result will b\[\frac{ 10! }{ 3!2! } \times \frac{ 5! }{ 2! }\]